OpenStudy (anonymous):

find the square root of (1-i)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf \sqrt{1-i} }\) like this ?

OpenStudy (solomonzelman):

you can't simplify this root.

OpenStudy (anonymous):

yeah there is an answer in the book ..

OpenStudy (solomonzelman):

The question here would be can you put whatever is in the square root as x^2, saying can you factor it in such a way that it's a perfect square. For example, \(\Large\color{blue}{ \bf \sqrt{2i} =\sqrt{(1+i)(1+i)}=\sqrt{(1+i)^2}=1+i }\) but your's 1-i is not a perfect square.

OpenStudy (asnaseer):

@Darkk_Reaper - do you know that any complex number can be written in the form \(re^{i\theta}\) ?

OpenStudy (anonymous):

no

OpenStudy (asnaseer):

what methods have you been taught for this kind of problem?

OpenStudy (solomonzelman):

that would probably be too complicated... to write it as \(\large\color{blue}{ \bf re^{iθ} }\) I usually either find a simple root, or leave it as it is.

OpenStudy (asnaseer):

e.g. do you know any complex number can also be written in the form \(r(\cos(\theta)+i\sin(\theta)))\)

OpenStudy (anonymous):

yeah

OpenStudy (asnaseer):

There is also a proof that shows that\[\cos(\theta)+i\sin(\theta)=e^{i\theta}\]If you know this then it simplifies this question

OpenStudy (asnaseer):

Otherwise you the methods shown here to solve this: http://www.mathcentre.ac.uk/resources/Engineering%20maths%20first%20aid%20kit/latexsource%20and%20diagrams/7_5.pdf

OpenStudy (asnaseer):

try it out - let me know if you get stuck and I'll try to explain it further

OpenStudy (anonymous):

THANX DUDE

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

yeah i have seen that , but how can i use those to solve this problem

OpenStudy (asnaseer):

let me formulate my thoughts on this as I would normally have solved this using \(re^{i\theta}\)

OpenStudy (asnaseer):

@Darkk_Reaper - do you know DeMoivre’s Theorem? i.e. \((a + ib)^n=r^n(\cos(n\theta)+i\sin(n\theta))\)

OpenStudy (anonymous):

SORRY